# The Causal Effect of the Two-For-One Strategy in the National Basketball Association

### v1

@@ 4 Results
In this section, we present the results of estimating the overall causal effect of the TFO strategy using two methods: IPW and propensity score matching. We then examine potential heterogeneity in treatment effects by employing the causal forest framework.
@@ 4.1 Average Treatment Effect
In our data set there are approximately twice as many treated observations as control observations each year. Therefore, when performing propensity score matching, it is more natural to estimate the average treatment effect on the control group, the ATC. For the sake of consistency, we estimate the ATC when using IPW as well. For teams that do not attempt a TFO in a certain situation, the ATC is a measure of how their average POD would have been different if they had attempted the TFO. This is a natural estimand for the goal of convincing teams that do not currently emphasize the TFO strategy to begin to do so. A positive ATC indicates an increase in the point differential, implying that the TFO strategy is beneficial. Conversely, if the ATC is not significantly different from zero or is negative, it suggests that the TFO strategy is ineffective.
The ATC is computed separately for each season, as well as for the combined dataset that includes both seasons. Estimating the ATC separately provides insight into the stability of the treatment effect across seasons, while combining the data allows for a larger sample size, leading to narrower confidence intervals. We report both the point estimates of ATC and the corresponding 95% confidence intervals.
We perform all calculations in R (R Core Team, 2024). To estimate the ATC using IPW, we employ the R package WeightIt to obtain the weights using covariate balancing propensity scores and the survey package to obtain point estimates and confidence interval. For propensity score matching, we obtain matches using the MatchIt package with 1:1 optimal pair matching without replacement, again based on covariate balancing propensity scores. We use the marginaleffects R package (Arel-Bundock et al., 2024) to obtain point estimates using g-computation as well as interval based on cluster-robust standard errors.
@@ 4.1.1 Checks for Assumptions
As mentioned in Sections 3.1 and 3.2, both IPW and matching rely on conditional exchangeability in order to create balanced groups. Figure 1 is a Love plot created in R for the combined sample using the cobalt package (Greifer, 2024a), which compares the standardized mean differences between the treatment and control groups of the various covariates before and after IPW and matching. In the original sample, five of the covariates have standardized differences that fall outside of the rule of thumb threshold of 0.1, including the time left variable with a value of .074. This indicates that the TFO-A and TFO-NA groups vary considerably, particularly in the amount of time left in the quarter. This is an expected feature of the data and highlights why causal methods are necessary.
Figure 1: Standardized mean differences (treatment-control) of the covariates before and after matching and IPW. The dotted line represents the rule of thumb threshold of 0.1.
Conversely, after weighting the observations, the standardized mean differences are very close to 0. This near perfect balance is in large part due to the use of covariate balancing propensity scores. Matching also creates more similar groups than in the original data set, although the gains are not nearly as pronounced as with IPW. The time left variable still has a standardized mean difference slightly above 0.1, although the rest are under the that threshold. In the combined data set there are 991 TFO-NA and 1986 TFO-A. The slightly worse balance for the matching method indicates that it was difficult to find quality matches for all of the TFO-NA given the relatively small number of TFO-A. Thus we should be slightly cautious when making conclusions from the matching method, while we can be more confident in the conclusions from IPW. Love plots for the individual seasons (not pictured) show a similar pattern to the combined data set.
@@
The positivity assumption can also be assessed after fitting the IPW and matching models. Typically this is done by looking at the distribution of the estimated propensity scores. When estimating the ATC, it is important that the support of the control group is roughly contained within the support of the treatment group. Figure 2 gives histograms of the propensity scores for both groups. Other than perhaps a couple of control observations with the smallest estimated propensity scores, there is complete overlap between the two distributions. Thus we can say that the positivity assumption is reasonable for the combined data set. Similar plots for the individual seasons (not pictured) reveal that the positivity assumption is reasonable in those analyses as well.
@@ 4.1.2 Treatment Effect Estimates
Method Season Average Treatment Effect
Estimate 95% Confidence Interval pvalue
IPW 18-19 0.55 (0.43, 0.66) $<0.001$
21-22 0.57 (0.45, 0.69) $<0.001$
Both 0.55 (0.47, 0.64) $<0.001$
Matching 18-19 0.60 (0.35, 0.85) $<0.001$
21-22 0.64 (0.39, 0.89) $<0.001$
Both 0.63 (0.46, 0.80) $<0.001$
@@
Table 2: ATC estimates based on IPW and matching for both the individual and combined seasons.
As shown in Table 2, the estimated ATC for both IPW and matching is consistently positive, with the 95% confidence intervals above 0 in all cases. For the IPW method, the ATC estimate for the combined dataset is 0.55 with a 95% confidence interval of (0.47, 0.64), while for the matching method, the combined ATC is slightly larger in magnitude at 0.63 with a confidence interval of (0.46, 0.80).
@@
While there are slight differences in the point estimates between the two methods and across individual seasons, the substantial overlap in confidence intervals suggests that these variations are not significant. Overall, the poistive ATC across both methods and seasons indicates that the TFO strategy consistently increases the points differential by the end of the period. On average, each time a team had a TFO-O and failed to convert it to a TFO-A, it cost itself slightly more than half a point of score differential. This may seem like a minor disadvantage, but NBA teams are often willing to spend effort and resources in pursuit of even relatively small edges.
@@ 4.2 Heterogeneous Treatment Effect
The average treatment effect provides valuable insight, indicating that the TFO strategy is beneficial on average. However, there is a need to explore how this effect varies across different values of the covariates. Understanding this heterogeneity can help identify specific scenarios where the TFO strategy should or should not be attempted, thereby informing team strategy.
To this end, we investigate treatment effect heterogeneity to address the following key questions. First, is there significant heterogeneity in the treatment effect? Second, which variables contribute to this heterogeneity? We employ the causal forest framework to answer these questions.
Before fitting the causal forest model, we augment the dataset with three additional variables: the difference in the maximum ratings between the two teams, the difference in the median ratings between the two teams, and the absolute value of the score margin. Initially, we run the causal forest model with all available variables. We then compute the variable importance measures and retain only those variables that cumulatively account for 95% of the total importance, eliminating three variables in the process. This step helps control the standard error.
As an initial check, we estimate the ATE using the causal forest. The estimated ATE is 0.61, with a standard error of 0.11, relatively close to the results obtained from IPW and matching methods.
This analysis sets the foundation for a deeper exploration of heterogeneity in treatment effects, guiding future strategy development.
@@ 4.2.1 Overall Heterogeneity
The first test we employ is the test calibration test, which evaluates the null hypothesis that there is no heterogeneity in the treatment effect. Table 3 shows the results of the test. The p-value of the differential forest prediction is 0.17, indicating that we do not have sufficient evidence to reject the null hypothesis. It is important to note that the standard error of this test is relatively high, which contributes to the high p-value and weakens our ability to detect heterogeneity, if present.
@@
Mean Forest Prediction 1.00511 0.17876 5.6227 $<0.001$
@@
The second test we conduct is based on the Rank Average Treatment Effect (RATE) (Yadlowsky et al., 2024). In this test, we define a prioritization rule and assess the impact of applying the treatment to only the top observations based on this rule. To implement this, we train the causal forest using 2018 data, predict the heterogeneous treatment effects for 2021 observations, and then use these predictions to form our prioritization rule. The RATE estimate is $0.035$ with a standard error of $0.112$. The standard error is estimated using bootstrap method (Efron, 1982) with 200 samples. Similar to the first test, this result suggests that there is insufficient evidence to reject the null hypothesis of no heterogeneity in the treatment effect. Figure 3 shows the overall TOC where we can see looking at the confidence band that 00 is in the confidence band.
Figure 3: Target Operating Curve generated using a causal forest model. This curve illustrates the estimates the treatment effect if only the top qof observations is given the treatment. The dashed line shows the 95% confidence bars.
In summary, both the test calibration and RATE results indicate that, based on our tests, we do not find strong evidence of treatment effect heterogeneity.
@@ 4.2.2 Heterogeneity Across Variables
In the previous section, we observed that at an aggregate level, the tests suggest minimal evidence for heterogeneity in the treatment effect. However, prior studies have demonstrated scenarios where overall homogeneity does not preclude the presence of heterogeneity across specific variables. To investigate this possibility, we examine heterogeneity across the following variables: difference in ratings of the two teams based on max rating, and difference in ratings of the two teams based on mean rating.
To evaluate potential heterogeneity, we employ the Rank Average Treatment Effect (RATE) framework. This method constructs prioritization rules based on the value of the selected variable and computes RATE estimates, reflecting the expectation that higher values of the variable should correspond to improved outcomes under treatment.
Table 4 presents the RATE estimates for the selected variables. The results indicate that the estimates are not statistically significant, providing insufficient evidence to conclude that treatment effect heterogeneity exists with respect to these variables. Thus, we conclude that there is no substantial evidence for heterogeneity in the treatment effect across the investigated variables.
@@ 5 Discussion
This study explores the effectiveness of the two-for-one (TFO) strategy in the NBA, converting a real-world sports strategy into a mathematical framework to enable causal inference. Using play-by-play data from the 2018-19 and 2021-22 NBA seasons, we analyze the impact of implementing this strategy on game outcomes by estimating the increase in the point differential a team that did not attempt a TFO could expect if it had attempted the TFO instead. Our findings indicate that attempting the TFO has a statistically significant positive effect, suggesting it can be advantageous for teams in real game scenarios.

### v3

@@ 4 Results
In this section, we present the results of estimating the overall causal effect of the TFO strategy using AIPW. We then examine potential heterogeneity in treatment effects by employing the causal forest framework. We perform all calculations in R (R Core Team, 2024). We implement the AIPW estimator using the CausalGAM package (Glynn and Quinn, 2017). We fit the propensity score model using a generalized additive model for a binomial outcome with a probit link function. We fit the outcome models using a generalized additive model for a conditionally Gaussian outcome with an identity link function. We use an empirical sandwich estimator of the sampling variance from Lunceford and Davidian (2004) to conduct inference.
@@ 4.1 Average Treatment Effect
We compute the ATE separately for each season, as well as for the combined dataset that includes both seasons, using the methods described in Section 3.2. Estimating the ATE separately provides insight into the stability of the treatment effect across seasons, while combining the data allows for a larger sample size, leading to narrower confidence intervals. For both the propensity score and outcome models we include all the covariates in Table 1. Additionally, for the combined data set we include season as a covariate.
@@ 4.1.1 Checks for Assumptions
As mentioned in Sections 3.1 and 3.2, AIPW relies on conditional exchangeability in order to create balanced groups. Figure 3 is a Love plot created in R for the combined sample using the cobalt package (Greifer, 2024), which compares the standardized mean differences between the treatment and control groups of the various covariates before and after AIPW. In the original sample, eight of the covariates have standardized differences that fall outside of the rule of thumb threshold of 0.05, including the time left variable with a value of 0.74. This indicates that the groups of TFO attempts and non-attempts vary considerably, particularly in the amount of time left in the period. This is an expected feature of the data and highlights why causal methods are necessary.
Figure 3: Standardized mean differences (treatment-control) of the covariates before and after AIPW. The dotted line represents the rule of thumb threshold of 0.05.
Conversely, after weighting the observations, the standardized mean differences are much closer to 0, with all covariates falling inside the threshold of 0.05. Love plots for the individual seasons (not pictured) show a similar pattern to the combined data set. Thus the weighting procedure has created groups that are comparable in terms of the measured confounders, providing evidence that the conditional exchangeability assumption is reasonable assuming the causal model represented by our DAG. In Section 4.1.3 we address the possibility of unmeasured confounders with a sensitivity analysis.
The positivity assumption can also be assessed after fitting the AIPW model. Here we do so by looking at the distribution of the estimated propensity scores. Figure 4 gives histograms of the propensity scores for both groups. Other than perhaps a couple of control observations with the smallest estimated propensity scores, there is complete overlap between the two distributions. Thus we can say that the positivity assumption is reasonable for the combined data set. Similar plots for the individual seasons (not pictured) reveal that the positivity assumption is reasonable in those analyses as well.
@@ 4.1.2 Treatment Effect Estimates
Season Average Treatment Effect
Estimate 95% Confidence Interval p-value
18-19 0.55 (0.31, 0.78) <0.001<0.001
21-22 0.77 (0.53, 1.00) <0.001<0.001
Both 0.66 (0.49, 0.83) <0.001<0.001
@@
Table 3: ATE estimates, 95% confidence intervals, and p-values for a test of a non-zero effect based on AIPW for both the individual and combined seasons.
@@
Table 3 gives the treatment effect estimates and 95% confidence intervals for both individual seasons and the combined data set. We estimate that if teams attempted all their TFO opportunities their average POD would be about 0.66 points higher than if they did not attempt any of their TFO opportunities, based on the combined data. A 95% confidence interval for this effect is (0.49, 0.83). The 2021-2022 ATE estimate of 0.77 is slightly higher than the 2018-2019 estimate of 0.55, but there is considerable overlap in their confidence intervals, suggesting that the variation across individual seasons may not be significant. Overall, the positive ATE across both seasons indicates that the TFO strategy consistently increases the points differential by the end of the period. The slightly more than half a point gain may seem like a minor advantage, but NBA teams are often willing to spend effort and resources in pursuit of even relatively small edges.
4.1.3 Sensitivity to Unmeasured Confounding
While we have utilized a reasonable set of covariates based on the available data, there exists the possibility of unmeasured confounding, a challenge to many causal analyses. It is natural to ask the question of how different our estimated treatment effects might be in the presence of such unmeasured confounding. Zhao et al. (2019) utilize a marginal sensitivity model to construct bootstrap confidence intervals for AIPW estimators based on a parameter $\Lambda$ that measures the degree of unmeasured confounding. Specifically, in the case of a single binary unmeasured confounder $U$, the odds of an individual with $U=1$ being in the treatment group is $\Lambda$ times higher than the odds of an individual with $U=0$ being in the treatment group. As a concrete example, suppose that immediately upon his team gaining an opportunity, a coach gives a recommendation to his players about whether or not to attempt the TFO. This could certainly have an effect on whether or not the players actually attempt the TFO. Suppose that situations in which the coach recommends the attempt have $\Lambda$ times higher odds to end up as an attempt than situations in which the coach does not recommend an attempt.
Figure 5 gives 95% confidence intervals for the ATE of a TFO attempt on the POD from the combined data set as a function of $\Lambda$ values ranging from 1.05 to 1.5, calculated using the bootsens R package (Zhao, 2018). Note that the intervals are entirely above 0 for values of $\Lambda$ up to 1.35, but contain 0 for $\Lambda\geq 1.4$ Using our example, this would mean that even if the odds of attempting a TFO increase by up to 35% when a coach recommends an attempt, the treatment effect would still be statistically significant for a two-sided test at a significance level of $\alpha=.05$. However, if those odds increase by more than 40%, the effect would no longer be statistically significant. For the 2018-2019 and 2021-2022 seasons, respectively, the treatment effect would still be statistically significant for $\Lambda$ values of up to 1.2 and 1.4, respectively.
Figure 5: 95% confidence intervals for the ATE given varying degrees of unmeasured confounding indicated by $\Lambda$.
For a more concrete example, consider the TFO attempt in the second period of the game between the Warriors and the Suns mentioned in Section 2.3. Our model estimated the propensity score for that TFO opportunity at 0.6654, which would imply the odds of attempting the TFO were around 2. If the true propensity score, given an unmeasured confounder, was 0.7, the odds would be 2.33 which would be 1.17 times higher than our estimate. However, if the true propensity score, given an unmeasured confounder, was 0.75, the odds would be 3 which would be 1.5 times higher than our estimate. In both individual seasons as well as the combined season, an overall $\Lambda=1.17$ would not be enough to change our conclusion, but an overall $\Lambda=2$ would.
Thus our results are somewhat sensitive to unmeasured confounding, particularly for the 2018-2019 season. However, in this regard having a treatment definition that relies on actions, rather than intent, may be beneficial, as it may lessen the strength of any potential confounding relationship. In our hypothetical example, due to the complexities of an NBA possession and the general autonomy of NBA players, the percentage of attempts out of all TFO opportunities may not be substantially different based on the coach’s recommendation or other unmeasured confounders.
4.1.4 Sensitivity to Treatment Definition
As mentioned in Section 2.1, another potential concern with our analysis is the somewhat arbitrary nature of our treatment definition. We chose this definition based on the structure of NBA games, before looking at any data. However, we allow that other analysts or NBA personnel might argue for slightly different cutoffs in the TFO definition. To investigate the robustness of our treatment effect estimate to these differences, we estimate the ATE of the TFO strategy for the combined data set, along with 95% confidence intervals, for a variety of other possible cutoffs. These results are presented in Figure 6. The horizontal axis represents the beginning and end of the TFO opportunity window, while the panels represent the cutoff for a TFO attempt. Thus the left-most category, 38-30, in the bottom panel represents a definition in which a team needs to gain possession with between 38 and 30 seconds left in the period to gain a TFO opportunity, and must take a shot or get fouled with at least 29 seconds left in the period for the opportunity to be considered an attempt. The vertical axis is the estimated ATE. The red circle represents the original definition with cutoffs of 43, 35, and 28 seconds, respectively.
Figure 6: Treatment effect estimates and 95% confidence intervals for alternative treatment definitions.
The treatment effect estimate using our original definition is the largest among the combinations of cutoffs we tested. However, there is considerable overlap between the original confidence interval and those with cutoffs fairly similar to our original definition. Note that definitions represented by the extreme left and right sides of the figure are, in our opinion, much less realistic and much less likely to be supported by other analysts or NBA personnel. For a broad swath of reasonable cutoffs, the estimated treatment effect is between 0.4 and 0.6. Thus while the treatment effect may be slightly higher with our definition than with others, the practical impact on policy may be minimal.
4.1.5 Assessing Fourth Period Data Exclusion
In Section 2.2, we motivated the exclusion of fourth period and overtime possessions due to end-of-game strategic differences. This subsection provides empirical evidence supporting that choice. Table 4 reports average treatment effect estimates by game period for the 2021–2022 season. The first three periods yield similar estimates, whereas the estimate for the fourth period and overtime is substantially attenuated and statistically indistinguishable from zero. This pattern suggests that end-of-game possessions follow meaningfully different strategic dynamics than those in earlier periods, justifying their exclusion from our main analysis.
Period Average Treatment Effect
Estimate 95% Confidence Interval p-value
1 0.81 (0.39, 1.23) <0.001<0.001
2 0.76 (0.31, 1.20) <0.001<0.001
3 0.60 (0.20, 1.00) $0.004$
4/OT 0.12 (-0.22, 0.46) $0.463$
Table 4: ATE estimates, 95% confidence intervals, and p-values for a test of a non-zero effect based on AIPW for the 2021-2022 season. These effects are conditional on the period, with the fourth period and overtime periods combined.
@@ 4.2 Heterogeneous Treatment Effect
The average treatment effect indicates that the TFO strategy is beneficial on average, but its value may vary across game contexts. To explore when the strategy may be more or less effective, we examine heterogeneous treatment effects (HTE) using causal forests.
We augment the dataset with three additional variables—the differences in the teams’ maximum and median ratings, and the absolute score margin—which capture relative team strength and game competitiveness. We first fit a causal forest using all variables, compute variable importance, and retain those accounting for 95% of total importance (excluding only period indicators and absolute score margin). As a check, the causal-forest ATE for the combined data set is 0.61 (SE 0.11), consistent with AIPW estimates.
@@ 4.2.1 Overall Heterogeneity
First, the test calibration procedure evaluates the null of no treatment effect heterogeneity. Once the causal forest is trained, heterogeneity is assessed via a Calibration Test (Chernozhukov et al., 2018), which regresses the outcome on the mean forest prediction (ATE) and the out-of-bag differential forest prediction (CATE). As shown in Table 5, the differential forest prediction yields a p-value of 0.17, providing no significant evidence against the null.
@@
Mean Forest Prediction 1.00511 0.17876 5.6227 <0.001<0.001
@@
As a complementary approach, we also compute the RATE statistic (Yadlowsky et al., 2025) RATE evaluates whether a prioritization rule based on predicted CATEs improves outcomes when treatment is selectively applied. We train the causal forest on 2018 data, predict CATEs for 2021, and then form the prioritization rule. The resulting RATE estimate is 0.0408 with a standard error of 0.1176, obtained via bootstrap (Efron, 1982) with 1000 samples. The confidence interval includes zero, indicating that the prioritization rule does not identify subgroups with reliably higher gains from treatment. This aligns with the test calibration results, reinforcing the conclusion that we lack strong statistical evidence of heterogeneity.
@@ 4.2.2 Heterogeneity Across Variables
Although global tests suggest limited evidence of heterogeneity, it is still possible that treatment effects vary systematically with respect to specific variables. Prior work (De Bosscher et al., 2009) highlights settings in which overall tests fail to detect heterogeneity, yet meaningful variation appears along particular covariates. Motivated by this, we examine two measures of relative team strength: the difference in maximum player ratings between teams and the difference in median player ratings. These variables are natural candidates for effect moderation, as differences in team quality could plausibly influence the payoff of attempting a TFO.
We again rely on the RATE framework, constructing prioritization rules based solely on each variable’s value. This allows us to test whether games featuring larger rating differences exhibit systematically different treatment effects. Table 6 summarizes the RATE estimates. Both are small in magnitude, with standard errors of comparable size, and neither is statistically significant. This suggests that even along these targeted axes—where we might reasonably expect strategic differences—the data do not provide evidence that the effectiveness of the TFO strategy changes meaningfully. Thus, we conclude that there is no substantial evidence for heterogeneity in the treatment effect across the investigated variables.
@@ 5 Discussion
This study explores the effectiveness of the two-for-one strategy in the NBA, converting a real-world sports strategy into a mathematical framework to enable causal inference. Using play-by-play data from the 2018-19 and 2021-22 NBA seasons, we analyze the impact of implementing this strategy on game outcomes by estimating the average increase in the point differential if all teams presented with a TFO opportunity attempt them versus if they do not attempt them. Our findings indicate that attempting the TFO has a statistically significant positive effect, suggesting it can be advantageous for teams in real game scenarios.